# Multi-scale Latent Point Consistency Models for 3D Shape Generation

### v2

@@ IV-D Multi-Class 3D Shape Generation
Following [3], we also jointly trained the multi-scale latent diffusion model and the MLPCM model across 13 different categories (airplane, chair, car, lamp, table, sofa, cabinet, bench, phone, watercraft, speaker, display, ship, rifle). Due to the highly complex and multi-modal data distribution, jointly training a model is challenging.
@@ IV-E Comparison with Nearest Neighbors and Voxel-guided Shape Synthesis
To supplement the conditional generation experiment to observe the generalization diversity of the model, We add voxel guided generation results as shown in Fig. 6. We begin by voxelizing the training set and fine-tuning our encoder networks to learn representations that can be accurately decoded back into the original shapes. For voxelized inputs processed with our point-cloud networks, we further sample surface points from the voxels to serve as input data.
@@ IV-F Sampling Time
The proposed multi-scale latent diffusion model synthesizes shapes using 1000 steps of DDPM, while MLPCM synthesizes them using only 1-4 steps, generating high-quality shapes in under 0.5 seconds. This makes real-time interactive applications feasible. In Table IV, we compare the sampling time and quality of generating a point cloud sample (2048 points) from noise using different DDPM, DDIM [31] sampling methods, and MLPCM. When using $\leq 10$ steps, DDIM’s performance degrades significantly, whereas MLPCM can produce visually high-quality shapes in just a few steps or even one step.
@@ IV-G Ablation Study
In this section, we conduct the ablation study to investigate the effectiveness of the individual design components introduced in the multi-scale latent diffusion model, including the multi-scale latents, 3D Spatial Attention Mechanism, and the additional dimensions of the latent points.
@@
Ablation on different backbones. We conduct an ablation study to evaluate different neural network architectures for processing point clouds in the encoder, decoder, and diffusion prior. The results, presented in Table VII, focus on the car category using the teacher model, consistent with other ablation studies. We test four popular backbone architectures: Point-Voxel CNN (PVCNN) [68], Dynamic Graph CNN (DGCNN) [21], PointTransformer [80], and PointTransformerV2 [81]. Among these, PVCNN demonstrates the best overall performance, making it our architecture of choice.
Ablation on 3D Positional Encoding. We conduct an ablation study to evaluate different 3D positional encoding methods. In the Voxel Positional Encoding (VPE) approach, we applied frequency-based sine-cosine 3D positional embeddings to the point-voxel branch in the SA module, following the design of DiT-3D [17]. For Contextual Relative Positional Encoding (cRPE), inspired by the Stratified Transformer [82], we enhanced the traditional positional encoding in the attention mechanism by adding a vector representing the positional offset between Query and Key during each attention computation. The experiments were performed on the car category, with results summarized in Table VIII. The findings show that omitting 3D positional encoding leads to significantly poorer generation results. While VPE and cRPE achieve comparable outcomes, we selected VPE as our 3D positional encoding method for its simplicity and effectiveness.
@@ V Conclusion
In this paper, we propose the Multi-Scale Latent Point Consistency Model (MLPCM) to address the task of efficient point-cloud-based 3D shape generation. We first construct a multi-scale latent diffusion model, which includes a point encoder based on multi-scale voxel convolutions, a multi-scale denoising diffusion prior with 3D spatial attention, and a voxel-convolution-based decoder. The diffusion prior model, which integrates multi-scale information, effectively captures both local and global geometric features of 3D objects, while the 3D spatial attention mechanism helps the model better capture spatial information and feature correlations. Moreover, MLPCM leverages consistency distillation to compress the prior into a one-step generator. On the widely used ShapeNet and ShapeNet-Vol datasets, the proposed multi-scale latent diffusion model achieves state-of-the-art performance, and MLPCM achieves a 100x speedup during sampling while surpassing the state-of-the-art diffusion models in terms of shape quality and diversity. In the future, we are interested in extending MLPCM to other diverse 3D representations such as meshes.

### v3

@@ IV-D Multi-Class 3D Shape Generation
Following [zeng2022lion], We also perform joint training of the multi-scale latent diffusion model and MLPCM over 13 ShapeNet categories, including airplane, chair, car, lamp, table, sofa, cabinet, bench, phone, watercraft, speaker, display, ship, and rifle. Due to the highly complex and multi-modal data distribution, jointly training a model is challenging.
@@ IV-E Comparison with Nearest Neighbors and Voxel-guided Shape Synthesis
To supplement the conditional generation experiment to observe the generalization diversity of the model, We add voxel guided generation results as shown in Fig. 6. We begin by voxelizing the training set and fine-tuning our encoder networks to learn representations that can be accurately decoded back into the original shapes. For voxelized shapes handled by our point-cloud networks, we additionally extract surface samples from the voxel grids and use these points as the network inputs.
@@ IV-F Sampling Time
The proposed multi-scale latent diffusion model synthesizes shapes using 1000 steps of DDPM, while MLPCM synthesizes them using only 1-4 steps, generating high-quality shapes in under 0.5 seconds. This makes real-time interactive applications feasible. In Table IV, we compare the sampling time and quality of generating a point cloud sample (2048 points) from noise using different DDPM, DDIM [song2020denoising] sampling methods, and MLPCM. When using $\leq 10$ steps, DDIM’s performance degrades significantly, whereas MLPCM can produce visually high-quality shapes in just a few steps or even one step.
@@ IV-G Ablation Study
In this section, we conduct an comprehensive ablation study to disentangle the effects of the main components in our multi-scale latent diffusion model. Specifically, we examine the impact of introducing multi-scale latent representations to capture both coarse global geometry and fine-grained local details, incorporating a 3D spatial attention mechanism to enable long-range dependency modeling and improved part-to-part coordination in 3D space, and augmenting the latent point representation with additional feature dimensions to increase expressiveness beyond pure coordinates. For each component, we report controlled comparisons under the same training protocol and evaluation metrics, and analyze how these choices affect generation fidelity, structural coherence, and sample diversity.
@@
Ablation on different backbones. We conduct an ablation study to evaluate different neural network architectures for processing point clouds in the encoder, decoder, and diffusion prior. The results, presented in Table VII, focus on the car category using the teacher model, consistent with other ablation studies. We test four popular backbone architectures: Point-Voxel CNN (PVCNN) [liu2019point], Dynamic Graph CNN (DGCNN) [wang2019dynamic], PointTransformer [zhao2021point], and PointTransformerV2 [wu2022point]. Among these, PVCNN demonstrates the best overall performance, making it our architecture of choice.
Ablation on 3D Positional Encoding. We conduct an ablation study to evaluate different 3D positional encoding methods. In the Voxel Positional Encoding (VPE) approach, we applied frequency-based sine-cosine 3D positional embeddings to the point-voxel branch in the SA module, following the design of DiT-3D [mo2023dit]. For Contextual Relative Positional Encoding (cRPE), inspired by the Stratified Transformer [lai2022stratified], we enhanced the traditional positional encoding in the attention mechanism by adding a vector representing the positional offset between Query and Key during each attention computation. The experiments were performed on the car category, with results summarized in Table VIII. The findings show that omitting 3D positional encoding leads to significantly poorer generation results. While VPE and cRPE achieve comparable outcomes, we selected VPE as our 3D positional encoding method for its simplicity and effectiveness.
@@ V Conclusion
[…]